OpenStudy (anonymous):

sove the equation 40x=4x2

OpenStudy (anonymous):

x*2 = 20 x= 2 Sqrt 5

OpenStudy (anonymous):

step by step please

OpenStudy (anonymous):

So, 40x=4x2, right? Well, you just do 4x2 which is 8. Then you have 40x=8. So then you do 40/8 and get 5. Does this help?

OpenStudy (anonymous):

its not 4x2 its 4squared

OpenStudy (anonymous):

Oh! Sorry. Then I don't know how to help you, sorry!

OpenStudy (anonymous):

lol it sok:)

OpenStudy (anonymous):

$$40x=4x^2$$ $$4x^2-40x=0$$ $$x^2-10x=0$$ $$x(x-10)=0$$ {0,10}